Can someone help me with these problems im stuck on? Thank you!

1)

Problem 12.88 Part A 0 g of aluminum at 200 °C and 20 g of copper are dropped into 43 cm3 of ethyl alcohol at 15 °C. The temperature quickly comes to 24 °C What was the initial temperature of the copper? Express your answer using two significant figures.

2)
Answer #1

here,

mass of alumunium , ma = 10 g

mass of copper, mc = 20 g

volume of ethyle alchocal , v = 43 cm^3

density of ethyle alchocal = 0.789 g/cm^3

mass of ethyle alchocal , me = 33.93 g

initial temprature of aluminium , Ta = 200 degree C

initial temprature of ethyl alchoal , Te = 15 degree C

specific heat of ethyle alchocal , Ce = 2.4 J/gm K

specific heat of alumunium, Ca = 0.9 J/gm K

specific heat of cooper , Cc = 0.386 J/gm K

let the initial temprature of copper be T

heat lost by aluminum + heat loss by cooper = heat gained by ethyl alchoal

ma * Ca * (Ta - 24 ) + mc * Cc * (24 - Tc) = me * Ce * (24 - Te)


10 * 0.9 * (200 - 24 ) + 20 * 0.386 * (24 - Tc) = 33.93 * 2.4 * (24 - 15)

Tc = 134.25 C

the initial temprature of copper is 134.25 degree C